# A proximal policy optimization based intelligent home solar management

### v1

@@ Abstract
This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible
@@ I Introduction
The goal of the intelligent agent is to manage solar energy and sell it based on the observation feed into a model. In this regard, we use Proximal Policy Optimization (PPO) and Sparse Mixture of Experts (MOE) for our time series forecasting algorithm for testing. Sparse Mixture of Experts is a method we use in our time series forecasting to allow larger models with less overfitting to get better generalizations on less data [1]. PPO is a training method for reinforcement learning that is trained for maximizing objectives in complex environments. PPO has shown to achieve state of the art in many tasks  [2].
@@
Our task is to perform smart home solar energy management. This problem requires the model to be both directly or indirectly capable of forecasting prices, and able to predict the most optimal time to accumulate rewards over time. In the literature, few works can be found that focus on cost minimization between the solar energy generated and the electricity demand from the different kinds of appliances  [3] [4]. Lee et al’s work combines both a supervised model to predict key variables like temperature to optimize the policy instead of just a simple AI agent [5]. This comes with the benefit of providing extra information to the AI agent. There are also other approaches that use optimization engines instead of reinforcement learning for management that show promising results over conventional methods [6]. There are other methods that show that applying game theory can help reduce the uncertainty of price changes in a dynamic market [7] and minimize the costs of energy utilization [8, 9]. Unlike these similar works, our problem is assuming no power usage is counted into the mix and it doesn’t focus on the distribution of energy to appliances. Our goal is to explicitly maximize the total accumulated profits from selling energy back to the grid. We feel that focusing on this allows us to apply our findings to other domains like natural language processing and managing financial markets.
@@
Most reinforcement algorithms like PPO are designed for a Markov decision process, which means that each state must not be reliant on the previous states. The Markov decision process with no adjustments presents limitations of a problem that does rely on previous states. There are methods to try to get around this like this work’s attempt to using Monte Carlo Tree Search policy trained using PPO [10]. The strength behind this approach is that it is able to plan ahead its actions and facilitate exploration in a stable way. This also gets more out of the decision process in the value estimation function. The main problem with using Monte Carlo Tree Search is the increasing memory requirements of storing each state in the tree. For every day there are two possible actions buying and selling. For every day there are two actions holding or selling. Therefore for each day $n$, there are $2^{n}$, where $n$ is days, possibilities. This makes it not feasible in our use case to use this method. Other methods try to optimize by using behavior cloning through reinforcement learning [11] to get around the Markov decision process, but our method does not rely on behavior cloning to find optimal rewards.
@@ II-B Generalized Advantage Estimation ( GAE)
We will use Generalized Advantage Estimation ( GAE)[13] for our advantage estimation function. In order to speed up training we attempted to vectored this and was able to get quicker training times. Our advantage function is modeled as follows.
@@ II-E Reccurent Rewards
By putting the rewards into the state, it allows the model to change its actions based on the previous state similar to how a recurrent neural network like LSTM stores previous information in its hidden state[14].

### v2

# biblio.bib

@@ Abstract
©2024 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works.
@@ I Introduction
The goal of the intelligent agent is to manage solar energy and sell it based on the observation feed into a model. In this regard, we use Proximal Policy Optimization (PPO) and Sparse Mixture of Experts (MOE) for our time series forecasting algorithm for testing. Sparse Mixture of Experts is a method we use in our time series forecasting to allow larger models with less overfitting to get better generalizations on less data [DBLP:journals/corr/ShazeerMMDLHD17]. PPO is a training method for reinforcement learning that is trained for maximizing objectives in complex environments. PPO has shown to achieve state of the art in many tasks  [DBLP:journals/corr/SchulmanWDRK17].
@@
Our task is to perform smart home solar energy management. This problem requires the model to be both directly or indirectly capable of forecasting prices, and able to predict the most optimal time to accumulate rewards over time. In the literature, few works can be found that focus on cost minimization between the solar energy generated and the electricity demand from the different kinds of appliances  [10066193] [HUY2023117340]. Lee et al’s work combines both a supervised model to predict key variables like temperature to optimize the policy instead of just a simple AI agent [s19183937]. This comes with the benefit of providing extra information to the AI agent. There are also other approaches that use optimization engines instead of reinforcement learning for management that show promising results over conventional methods [6599649]. There are other methods that show that applying game theory can help reduce the uncertainty of price changes in a dynamic market [8031035] and minimize the costs of energy utilization [6835827, 5628271]. Unlike these similar works, our problem is assuming no power usage is counted into the mix and it doesn’t focus on the distribution of energy to appliances. Our goal is to explicitly maximize the total accumulated profits from selling energy back to the grid. We feel that focusing on this allows us to apply our findings to other domains like natural language processing and managing financial markets.
@@
Most reinforcement algorithms like PPO are designed for a Markov decision process, which means that each state must not be reliant on the previous states. The Markov decision process with no adjustments presents limitations of a problem that does rely on previous states. There are methods to try to get around this like this work’s attempt to using Monte Carlo Tree Search policy trained using PPO [liu2023dont]. The strength behind this approach is that it is able to plan ahead its actions and facilitate exploration in a stable way. This also gets more out of the decision process in the value estimation function. The main problem with using Monte Carlo Tree Search is the increasing memory requirements of storing each state in the tree. For every day there are two possible actions buying and selling. For every day there are two actions holding or selling. Therefore for each day $n$, there are $2^{n}$, where $n$ is days, possibilities. This makes it not feasible in our use case to use this method. Other methods try to optimize by using behavior cloning through reinforcement learning [aruoss2024grandmasterlevel] to get around the Markov decision process, but our method does not rely on behavior cloning to find optimal rewards.
@@ II-B Generalized Advantage Estimation ( GAE)
We will use Generalized Advantage Estimation ( GAE)[schonlau_random_2020] for our advantage estimation function. In order to speed up training we attempted to vectored this and was able to get quicker training times. Our advantage function is modeled as follows.
@@ II-E Reccurent Rewards
By putting the rewards into the state, it allows the model to change its actions based on the previous state similar to how a recurrent neural network like LSTM stores previous information in its hidden state[10.1162/neco.1997.9.8.1735].